June 11, 2009 Calorimetry

From GlueXWiki
Jump to: navigation, search

Teleconference Time: 11:00 a.m. EST

  • EVO session
  • Phone:
    • +1-800-377-8846 : US
    • +1-888-276-7715 : Canada
    • +1-302-709-8424 : International
    • then enter participant code: 39527048# (remember the "#")

Items for followup from previous meeting(s)

  1. Resolution studies
    • Study impact of inactive material at the end of BCAL on acceptance.
  2. Documentation on fiber first-article
  3. Review calibration procedures
  4. Will: Noted from CLAS experience both oxides and carbonates can accumulate on lead. One type is worse than the other (will send references to UofR)
    • Note: no written documentaion
  5. Elton will check on fiber delivery and comments from Kuraray on material for storage of fibers.

Documents to Review

Tentative Agenda

  1. Announcements
  2. Review of Action Items
  3. Calibration
  4. FCAL Update
  5. BCAL Fabrication
  6. BCAL Readout
    • Update from USM (Will)
      • Report from Alam/Sergey: Measurements of the effect of cooling on SensL SiPMs
    • VERY Budget comparisons (Elton)
    • Discussion of Bcal Readout Review (July 21)
      • Stefano Miscetti can participate in the review via video. A test of ESNET connection yesterday was very successful.
      • Other reviewers?
      • Agenda
  7. Simulations


Attendees: USM: Alam, Will, Sergei, JLab: Elton, Beni, Fernando, Eugene, Sasha, Ashot; IU: Mat, Ryan, Claire, Mijahlo

  1. Review of Action Items
    • Will reported that he checked on CLAS documentation on lead contamination and found none, but oxides and carbonates should be investigated.
  2. Calibration
    • No report - need to schedule a time for input from Athens.
  3. FCAL Update
    1. Mou signed and moving up chain at IU. Should be returned to JLab next week.
    2. Prototype (64 Pb glass blocks) has 50 operational pmt bases.
      • LED pulsing system with plexiglass sheet in front
      • CAMAC daq system for readout is operation. Goals are to test pmt operation and dividers, look for cosmics
    3. Final B-shield sent to Tim for evaluation
      • Tim: received shield, will obtain a mu-metal shieldl from Amuneal and let magnet measurement group test combined shield in B-field
    4. PMT base as outdated micro chip; want to change vendors
      • summer student working with Claire
      • CAN net port to new vendor chips. Work wiht Elliott to make sure that slow control is handled in a compatible way with online
      • Target is to redesign base and have working prototype by fall.
    5. Rad Hard blocks, specifying number of blocks needed
      • Meeting with Paul Sokol to review options to use a medical linac (6-20 MeV electrons) to irradiate blocks
      • Purchase 2 blocks with reduced tolerances (cheaper) for radiation tests
      • Goal is to irradiate using linac and tune GEANT MC to reproduce damage.
      • Also investigating possibilities for using the medical linac into a storage ring to produce a 100 MeV extracted beam for tests.
    6. FADC work
      • Received FADC back from JLab
      • Jake has tested various systematic tests suggested at the collaboration meeting. No significant changes seen, limiting systematics to about 50 ps.
      • Have noticed problems with some channels in the FADC; will return to JLab for repair
      • Alternative timing algorithms are under investigation, including a linear interpolation scheme suggested by Mark.
      • Draft of FADC work is in progress.
  4. BCAL Fabrication
    1. Contract status. Elton reported that Regina has replied to the RFP and JLab is evaluating response.
    2. Construction prototype update (See contributed notes from Zisis following)
    3. Fiber QA: Media:AttenLen_JS001-8_Prelim.png
      • Measurements of production fibers have atten length of about 350 cm, with a sigma of 5%. (Satisfy specs)
    4. Fiber measurements (Andrei)
      • Discussion shows a kink in a fiber used for comparison measurements. Andrei showed the impact on Npe and lambda for pmt and Sipm measurements. New measurements reduce discrepancy in first measurements but does not eliminate them.
      • PMT used for maeasurements is the Hamamatsu R329-02.
      • Eugene: 30% non-linearity in ADC seems large can be checked very quickly.
      • Eugene asked about differences in fiber behavior for single fiber compared to fiber in matrix. Andrei felt that current conditions should not be too differnet than those inside the matrix.
  5. BCAL Readout
    1. Update from USM (Will)
      • Report from Alam/Sergey: Measurements of the effect of cooling on SensL SiPMs
      • Alam and Sengei are developing a 256 channel LV power supply for SiPMs. Using SensL 3x3 mm2 SiPM for evaluating the performance of the power supply.
      • Showed some dark current measurements at 6.4deg and 20.4 deg. Checking the measurements relative to the NIM article by Lightfoot
      • Eugene: asked if the bias voltage can be used to reduce gain by about a factor of 4. Seems possible, but with a loss of PDE efficiency. (This may limit flexibility in adjusting the gain to different operating conditions.)
    2. [VERY preliminary budget comparisons] (Elton)
      • Preliminary estimates of costs were presented. We still have some soft numbers for the cost of a) light guides b) temperature stabilization, both mechanical and electrical c) summing circuits for SiPMs. We need to firm these up.
    3. Discussion of Bcal Readout Review (July 21)
      • Stefano Miscetti can participate in the review via video. A test of ESNET connection yesterday was very successful.
      • Will has made travel plans and can attend the review on July 21.
      • Other reviewers?
      • Agenda
  6. Simulations

Notes submitted by Zisis:

  1. BCAL contract: George carried out revisions suggested by Teresa Danforth on how to report "burdened" costs.
  2. BCAL Construction
    • Installation of pressure lines and electrical has started.
    • Building of swager cubby hole delayed by 2-3 weeks. No impact on schedule in the absence of the second press.
    • Materials have been prepared for a last baby cal: this one 2m long, 12 cm wide, and with 20 layers. Dan machined an aluminum base plate for it, including the 1mm fiber guide groove. The first Pb sheet will be glued with industrial epoxy. This exercise is meant to accomplish:
      • Practice in gluing the first lead sheet to the Al plate with industrial epoxy.
      • Test the operation and use of the electropneumatic press.
      • Study of epoxy migration via staggered ram pressure over significant length.
      • First baby cal made using the press. Both ends will be machined and the resulting fiber pitch measured accurately. Use this information to finalize the roller gap for the Construction Prototype.
    • Lead for 3/4 of Construction Prototype has been prepared.
    • Fiber sorting will commence during the first week of July after fiber QA tests have been digested and vetted through the Cal WG. For the production modules QA testing and sorting will proceed in parallel, in an assembly line fashion. Now we are being extra cautious since this is our first large batch of fibers.
    • First four gallons of Bicron 600 glue are on their way. Assuming they arrive by the end of the month construction
    • All other supplies for construction are in hand.
    • Discussions and quotations from Ross and HydroMachine are well in hand.
  3. Fiber QA Tests
    • Fiber Care
      • Wavelength spectra of room and desktop lamp lights were collected for both our Detector Lab (LB127 - where QA takes places) and Construction Lab (LB113 - where construction takes place). Filters are different but both cut out pretty much everything below 450 nm. Fibers are exposed for less than a minute to fluorescent hallway lights during the extraction from their cardboard boxes. This is done when pulling them out in order to always keep them straight and not bend them in handling; LB127 is less than 8m long. Overhead room lights in LB127 are kept off, and only 1-2 small incandescent lamps are used for lighting; these also have filters in them.
      • The fibers are always handled with cotton gloves and kept in their poly bags as much as possible.
      • We plan to purchase and install temperature and humidity sensors in both labs.
    • After experimenting with various containers/trays, we decided to cut up use the Kuraray cardboard boxes instead of the rain gutters for single, tested fibers, for ease in handling. The gutters are large and can handle fiber bags. We kept some cardboard boxes intact in case we need them for shipping fibers to JLab or (God forbid) back to Kuraray. Shelving (>4m long) has been erected in both LB113 and LB127 to hold fiber trays.
    • Procedure: No fiber enters LB113 until its batch has been approved after the QA. This avoids any confusion on which fibers can be used in the construction.
    • Over 100 fibers (0.3% of total or 30% of to-be-tested ones) have been measured for attenuation length using a calibrated Hamatsu photodiode and a picoammeter. Preliminary results have been posted on today's Wiki page. Fibers meet specs. Only one is near (but still over) the 300 cm mark. No. of p.e. tests will start next week.
  4. ZP is back to working on fiber report and hopes to have a draft out during the week of June 22.